# On the equivalence of all notions of generalized derivations whose domain is a C-algebra

### v1

@@ 2. Automatic continuity of generalized derivations of the third type
The main goal of this section is to get a result on the automatic continuity of generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule. We begin with an example of a 3-tuple of maps behaving like a derivation in which one or two of the last two maps is not linear.
@@
Observe that $F,H$ is linear while $G$ is non-necessarily linear. A straightforward verification shows that $F(AB)=G(A)B+AH(B),$ for all $A,B\in\mathfrak{A},$ which means that $(F,G,H)$ behaves like a ternary derivation on $\mathfrak{A}$, and hence $F$ is a generalized derivation of the third type on $\mathcal{A}$.
Let $D$ be a derivation on $\mathcal{A}$. Consider the Banach algebra =$\mathcal{B}=\left\{\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):a\in\mathcal{A}\right\}$ and the $\mathcal{B}$-bimodule $\mathcal{M}=M_{2}(\mathcal{A})=\left\{\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):a,b,c,d\in\mathcal{A}\right\}$, with the obvious operations. Define $F,G,H:\mathcal{B}\to M_{2}(\mathcal{A})$, by $F\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&0\\ 0&0\end{array}\right),$ $G\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&f(a)\\ 0&0\end{array}\right),$ and $H\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&0\\ f(a)&0\end{array}\right).$ Clearly, $G$ and $H$ are non-necessarily linear, and is easy to see that $(F,G,H)$ behaves like a derivation.
It follows from the previous counterexample that if the triplet $(F,G,H)$ behaves like a derivation, the mappings $G$ and $H$ need not be, in general, linear. The existence of a unit element in $\mathcal{A}$ forces the linearity of these maps, up to an appropriate change on the module.
@@
Let $(F,G,H)$ behaves like a derivation from an algebra $\mathcal{A}$ to an $\mathcal{A}$-bimodule $\mathcal{M}$. Suppose $\mathcal{A}$ is unital with unit $\mathbf{1}$. Then the mappings $G\cdot\mathbf{1},\mathbf{1}\cdot H:\mathcal{A}\to\mathcal{M},$ $x\mapsto G(x)\mathbf{1}$, $x\mapsto\mathbf{1}H(x)$ are linear, and the $3$-tuple $(F,G\cdot\mathbf{1},\mathbf{1}\cdot H)$ behaves like a derivation. If $\mathcal{M}$ is unital, the mappings $G$ and $H$ are linear.
@@
The conclusion is clear from the identities $(G\cdot\mathbf{1})(a)=F(a)-aH(\mathbf{1}),$ $(\mathbf{1}\cdot H)(a)=F(a)-G(\mathbf{1})a$ ($a\in\mathcal{A}$) and the linearity of $F$. Observe that
@@
Let us recall some well known concepts. Let $\mathcal{A}$ be an algebra and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. The modular left annihilator of $\mathcal{A}$ is the set $lann(\mathcal{A})_{\mathcal{M}}:=\{m_{0}\in\mathcal{M}\ |\ m_{0}\mathcal{A}=\{0\}\}.$ Similarly, the modular right annihilator of $\mathcal{A}$ defined by $rann(\mathcal{A})_{\mathcal{M}}:=\{m_{0}\in\mathcal{M}\ |\ \mathcal{A}m_{0}=\{0\}\}.$ The modular annihilator of $\mathcal{A}$ is $ann(\mathcal{A})_{\mathcal{M}}:=rann(\mathcal{A})_{\mathcal{M}}\cap lann(\mathcal{A})_{\mathcal{M}}$. In particular, if $\mathcal{A}=\mathcal{M}$, then we reach the usual notions of left annihilator, right annihilator and annihilator of the algebra $\mathcal{A}$ which are denoted by $lann(\mathcal{A})$, $rann(\mathcal{A})$ and $ann(\mathcal{A})$, respectively.
Along this section, the left and right multiplication operators by an element $a$ in an associative algebra $\mathcal{A}$ will be denoted by $L_{a}$ and $R_{a},$ respectively. If $\mathcal{M}$ is an $\mathcal{A}$-bimodule, the mappings $L_{a}$ and $R_{a}$ will also stand for the corresponding left and right multiplication operators on $\mathcal{M}$.
@@
Let $\mathcal{A}$ be an algebra, let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. Suppose $(F,G,H)$ is a $3$-tuple of mappings from $\mathcal{A}$ to $\mathcal{M}$ behaving like a derivation. Then the following statements hold:
@@ 2. $(ii)$
    If $rann(\mathcal{A})_{\mathcal{M}}=\{0\}$ (respectively, $lann(\mathcal{A})_{\mathcal{M}}=\{0\}$), then the mapping $G$ (respectively, the mapping $H$) is linear.
@@
$(i)$ We show that the mapping $\Psi_{a}$ is a linear mapping for every $a\in\mathcal{A}$. Given $a,b,c\in\mathcal{A}$ and $\lambda\in\mathbb{C}$, we have
@@
Hence, $\Psi_{a}$ is a linear mapping for any $a\in\mathcal{A}$. Similarly, one can show that the mapping $\Gamma_{a}:\mathcal{A}\rightarrow\mathcal{M},$ $\Gamma_{a}(t)=G(t)a$ is a linear mapping.
@@
$H(\lambda(b+c))=\lambda H(b)+\lambda H(c)$
for all $b,c\in\mathcal{A}$ and all $\lambda\in\mathbb{C}$, and thus $H(a+b)=H(a)+H(b)$ for all $a,b\in\mathcal{A}$, $H(0)=0$ and $H(\lambda a)=\lambda H(a)$ for all $\lambda\in\mathbb{C}$. The other statement can be similarly obtained. ∎
@@
Let $\mathbb{Z}$ be the set of all integers. Set
$\mathcal{A}=\left\{\left[\begin{array}[]{cc}2n&0\\ 0&2n\end{array}\right]\ :\ n\in\mathbb{Z}\right\}.$
It is evident that $\mathcal{A}$ is a nonunital ring. Let
=$\mathcal{M}=\left\{\left[\begin{array}[]{cc}i&j\\ 0&k\end{array}\right]\ :\ i,j,k\in\mathbb{Z}\right\}.$
A straightforward verification shows that $lann(\mathcal{A})_{\mathcal{M}}=\{0\}=rann(\mathcal{A})_{\mathcal{M}}$.
If $\mathcal{A}$ is unital, $lann(\mathcal{A})=\{0\}=rann(\mathcal{A})$. Also, if $\mathcal{A}$ is a semiprime algebra, then it is clear that $lann(\mathcal{A})=rann(\mathcal{A})=\{0\}$. If $\mathcal{A}$ satisfies Condition (P) in [12], then it is routine to see that $lann(\mathcal{A})=rann(\mathcal{A})=\{0\}$, where an algebra or a ring $\mathcal{A}$ satisfies Condition (P) if $aa_{0}a=\{0\}$ for any $a\in\mathcal{A}$ implies that $a_{0}=0$.
The dual space, $\mathcal{A}^{*}$, of a Banach algebra $\mathcal{A}$ is a Banach $\mathcal{A}$-bimodule with the module operation given by $(\varphi a)(b)=\varphi(ab)$ and $(a\varphi)(b)=\varphi(ba)$, for all $a,b\in\mathcal{A}$, $\varphi\in\mathcal{A}^{*}.$ If $\mathcal{A}$ satisfies a kind of Cohen factorization property (i.e. for every $c\in\mathcal{A}$ there exists $a,b\in\mathcal{A}$ with $c=ab$), we have $lann_{\mathcal{A}^{*}}(\mathcal{A})=\{0\}=rann_{\mathcal{A}^{*}}(\mathcal{A})$. Recall that, by Cohen’s factorization theorem [19, Corollary 2.26], every Banach algebra with a bounded left approximate unit satisfies such a factorization property. To see the statement concerning modular annihilators, if $\varphi\in lann_{\mathcal{A}^{*}}(\mathcal{A}),$ we have $\varphi(ab)=0$ for all $a,b\in\mathcal{A}$, the factorization property implies that $\varphi=0$.
Our next proposition shows that continuity is the essential property to conclude that a generalized derivation of the third type from a C-algebra $\mathcal{A}$ into a general Banach $\mathcal{A}$-bimodule is a generalized derivation of the first and second type, improving in this way what we commented at the introduction.
@@
Let $\mathcal{A}$ be a C-algebra and let $F:\mathcal{A}\to\mathcal{M}$ be a continuous generalized derivation of the third type from $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule. Then $F$ is a generalized derivation of the first and second type.
@@
It is known that the product of $\mathcal{A}$ and the module products on $\mathcal{M}$ can be extended to a product on $\mathcal{A}^{**}$ and $\mathcal{A}^{**}$-bimodule operations on $\mathcal{M}^{**}$ via the first Arens extensions [10, Theorem 2.6.15$(iii)$], respectively –the second Arens extension is also valid. Since $\mathcal{A}$ is a C-algebra, its product is Arens regular, that is, the first and second Arens products on $\mathcal{A}^{**}$ coincide, and make the latter a von Neumann algebra (cf. [10, Corollary 3.2.37]). It is also known that the following properties holds: for each $a\in\mathcal{A}$, $\tilde{a}\in\mathcal{A}^{**}$, $x\in\mathcal{M}$ and $z\in\mathcal{M}^{**}$, the mappings $y\mapsto ay$, $y\mapsto y\tilde{a}$ (respectively, $b\mapsto xb$, $b\mapsto bz$) are weak continuous maps on $\mathcal{M}^{**}$ (respectively, from $\mathcal{A}^{**}$ to $\mathcal{M}^{**}$) [10, Proposition A.3.52]; if $(a_{\lambda})$ and $(x_{\mu})$ are nets in $A$ and $X$, respectively, such that $a_{\lambda}\to a\in A^{**}$ in the weak topology of $A^{**}$ and $x_{\mu}\to x\in X^{**}$ in the weak topology of $X^{**}$, then
$ax=\lim_{\lambda}\lim_{\mu}a_{\lambda}x_{\mu}\hbox{ and }xa=\lim_{\mu}\lim_{\lambda}x_{\mu}a_{\lambda}$ (2.2)
in the weak topology of $X^{**}$ (cf. [10, (2.6.26)]).
@@
and hence the uniform boundedness principle assures the existence of a positive $K_{1}$ satisfying $\|G(\cdot)b\|\leq K_{1}$ for all $b\in\mathcal{A}$ with $\|b\|\leq 1$. Similarly, there exists of a positive $K_{2}$ satisfying $\|aH(\cdot)\|\leq K_{2}$ for all $a\in\mathcal{A}$ with $\|a\|\leq 1$.
Let us take an approximate unit $(u_{j})_{j}$ in $\mathcal{A}$. If we fix an element $a\in\mathcal{A}$ the net $(au_{j})_{j}$ converges in norm to $a$, and hence by the continuity of $F$, $F(au_{j})$ tends to $F(a)$ in norm. It is also know that $(u_{j})_{j}\to\mathbf{1}$ in the weak-topology of $\mathcal{A}^{**}$, where $\mathbf{1}$ stands for the unit in $\mathcal{A}^{**}$. We therefore deduce from (2.2) that $(G(a)u_{j})_{j}\to G(a)\mathbf{1}$ in the weak-topology of $\mathcal{M}^{**}$. Thus, the identity $F(au_{j})=G(a)u_{j}+aH(u_{j})$ implies that
the net $(aH(u_{j}))_{j}\hbox{ converges in the weak${}^{*}$-topology of }\mathcal{M}^{**}$ (2.3)
@@
It is easy to check that the mapping $R:\mathcal{A}\to\mathcal{M}^{**},$ $a\mapsto R(a)$ is linear. Moreover, since by the properties shown above $\|aH(u_{j})\|\leq K_{2}\|a\|$ for all $j$, we obtain that $\|R(a)\|\leq K_{2}\|a\|$ for all $a\in\mathcal{A}$, and hence $R$ is continuous.
@@
$R(ab)=w^{*}\hbox{-}\lim_{j}abH(u_{j})=aw^{*}\hbox{-}\lim_{j}bH(u_{j})=aR(b),$
which guarantees that $R$ is a right multiplier from $\mathcal{A}$ to $\mathcal{M}^{**}$. Furthermore, $\|H(u_{j})\|\leq K_{2}$ for all $j$, and hence, by the weak-compactness of the closed unit ball of $\mathcal{M}^{**}$, there exists a subnet, denoted again by $(u_{j})_{j}$, such that $\lim_{j}H(u_{j})=\xi\in\mathcal{M}^{**}$ in the weak-topology of $\mathcal{M}^{**}$. Since $R(au_{j})=aR(u_{j})$ for all $j$, and $(au_{j})_{j}\to a$ in norm, the continuity properties of the module operations on $\mathcal{M}^{**}$ and $R$ give $R(a)=a\xi$ for all $a\in\mathcal{A}$.
@@
We shall next state two classical arguments in results on automatic continuity which are considered here under a more general point of view.
@@
Let $Z$ be a closed subspace of a Banach space $X$ such that $X/Z$ is finite dimensional, and let $Y$ be a normed space. Suppose $F:X\to Y$ is a linear mapping whose restriction to $Z$ is continuous. Then $F$ is continuous.
@@
Since the quotient $X/Z$ is finite dimensional, the subspace $Z$ is topologically complemented in $X$, that is, there exists a continuous linear projection $P:X\to X$ whose image is $Z$ and $Z^{\prime}=(Id-P)(X)$ is finite dimensional (cf. [39, Lemma 2.21]). Clearly, by the finite dimensionality of $Z^{\prime}$, $F|_{Z^{\prime}}:Z^{\prime}\to Y$ is continuous. Since $F|_{Z}$ is continuous by hypothesis, and $F(x)=F(P(x))+F((Id-P)(x))=F|_{Z}(x)+F|_{Z^{\prime}}(x)$ for all $x\in X$, the conclusion is clear. ∎
@@
The desired conclusion is clear when $\mathcal{A}$ is unital, since in that case $F(a)=FL_{\mathbf{1}}(a)$. In the general case, we observe that the bilinear mapping $(a,b)\mapsto V(a,b):=F(ab)$ is separately continuous by hypothesis, so by the uniform boundedness principle, $V$ is jointly continuous. Hence, there exists a positive $K$ such that $\|F(ab)\|\leq K\ \|a\|\ \|b\|$ for all $a,b\in\mathcal{A}$. For each positive element $c\in\mathcal{A}$ with $\|c\|\leq 1$ there exists a positive $d$ with $d^{2}=c$, $\|d\|^{2}=\|d^{2}\|=\|c\|\leq 1$. Therefore $\|F(c)\|=\|F(d^{2})\|\leq K\|d\|^{2}=K\|c\|\leq K$. In particular, $F$ is bounded on the closed unit ball of $\mathcal{A}$. ∎
A cornerstone result in the theory of C-algebras, obtained by Cuntz in [9], asserts that a semi-norm $p$ on a C-algebra $\mathcal{A}$ which is bounded on each commutative self-adjoint subalgebra of $\mathcal{A}$, is bounded on the whole of $\mathcal{A}$. Cuntz’ theorem has been employed in results on automatic continuity of derivations, for example, in [37], Russo and the second author of this note apply it to prove that every Jordan derivation from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule is continuous and an associative derivation. More recently, An and He employed Cuntz theorem to prove that if $n\neq m$, then zero is the only $(m,n)$-Jordan derivation from a C-algebra into a Banach bimodule (cf. [3]). We give another application next by adapting the arguments in the proof of Theorem 2.8.
We are now ready to prove one of the main results of this section, which confirm that every generalized derivation of the third type from a general C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule is continuous.
@@
Let $\mathcal{A}$ be a C-algebra, let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule, and let $F$ be a generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$. Then $F$ is continuous. Consequently, every generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the second type.

### v2

@@ 2. Automatic continuity of generalized derivations of the third type
The main goal of this section is to establish a result assuring the automatic continuity of every generalized derivation of the third type from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule. We begin with an example of a $3$-tuple of maps behaving like a derivation in which at least one of the last two maps in the triplet is not linear.
@@
Observe that $F$ and $H$ are linear while $G$ is non-necessarily linear. A straightforward verification shows that $F(AB)=G(A)B+AH(B),$ for all $A,B\in\mathfrak{A},$ which means that $(F,G,H)$ behaves like a ternary derivation on $\mathfrak{A}$, and hence $F$ is a generalized derivation of the third type on $\mathcal{A}$.
Let $D$ be a derivation on $\mathcal{A}$. Consider the Banach algebra =$\mathcal{B}=\left\{\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):a\in\mathcal{A}\right\}$ and the $\mathcal{B}$-bimodule $\mathcal{M}=M_{2}(\mathcal{A})=\left\{\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):a,b,c,d\in\mathcal{A}\right\}$, with the natural operations. Define $F,G,H:\mathcal{B}\to M_{2}(\mathcal{A})$, by $F\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&0\\ 0&0\end{array}\right),$ $G\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&f(a)\\ 0&0\end{array}\right),$ and $H\left(\begin{array}[]{cc}a&0\\ 0&0\end{array}\right):=\left(\begin{array}[]{cc}D(a)&0\\ f(a)&0\end{array}\right).$ Clearly, $G$ and $H$ are non-necessarily linear, and it is easy to see that $(F,G,H)$ behaves like a derivation.
It follows from the previous counterexample that if the triplet $(F,G,H)$ behaves like a derivation, the mappings $G$ and $H$ need not be, in general, linear. However, if $\mathcal{A}$ is unital, we can replace them by linear mappings.
@@
Let $\mathcal{A}$ be an algebra, and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. Suppose $(F,G,H)$ is a $3$-tuple of mappings from $\mathcal{A}$ to $\mathcal{M}$ behaving like a derivation. Assume that $\mathcal{A}$ is unital with unit $\mathbf{1}$. Then the mappings $G\cdot\mathbf{1},\mathbf{1}\cdot H:\mathcal{A}\to\mathcal{M},$ $x\mapsto G(x)\mathbf{1}$, $x\mapsto\mathbf{1}H(x)$ are linear, and the $3$-tuple $(F,G\cdot\mathbf{1},\mathbf{1}\cdot H)$ behaves like a derivation. If $\mathcal{M}$ is unital, the mappings $G$ and $H$ are linear.
@@
The conclusion is clear from the linearity of $F$, the identities $(G\cdot\mathbf{1})(a)=F(a)-aH(\mathbf{1}),$ $(\mathbf{1}\cdot H)(a)=F(a)-G(\mathbf{1})a$ ($a\in\mathcal{A}$), and
@@
Let us recall some well-known concepts. Let $\mathcal{A}$ be an algebra and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. The modular left annihilator of $\mathcal{A}$ is the set $lann(\mathcal{A})_{\mathcal{M}}:=\{m_{0}\in\mathcal{M}\ |\ m_{0}\mathcal{A}=\{0\}\}.$ Similarly, the modular right annihilator of $\mathcal{A}$ defined by $rann(\mathcal{A})_{\mathcal{M}}:=\{m_{0}\in\mathcal{M}\ |\ \mathcal{A}m_{0}=\{0\}\}.$ The modular annihilator of $\mathcal{A}$ is $ann(\mathcal{A})_{\mathcal{M}}:=rann(\mathcal{A})_{\mathcal{M}}\cap lann(\mathcal{A})_{\mathcal{M}}$. In particular, if $\mathcal{A}=\mathcal{M}$, we find the usual notions of left annihilator, right annihilator, and annihilator of the algebra $\mathcal{A},$ which are denoted by $lann(\mathcal{A})$, $rann(\mathcal{A}),$ and $ann(\mathcal{A})$, respectively.
Along this section, the left and right multiplication operators by an element $a$ in an associative algebra $\mathcal{A}$ will be denoted by $L_{a}$ and $R_{a},$ respectively. If $\mathcal{M}$ is an $\mathcal{A}$-bimodule, the mappings $L_{a}$ and $R_{a}$ will also stand for the corresponding left and right multiplication operators on $\mathcal{M},$ respectively.
@@
Let $\mathcal{A}$ be an algebra, and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. Suppose $(F,G,H)$ is a $3$-tuple of mappings from $\mathcal{A}$ to $\mathcal{M}$ behaving like a derivation. Then the following statements hold:
@@ 2. $(ii)$
    If $rann(\mathcal{A})_{\mathcal{M}}=\{0\}$ (respectively, $lann(\mathcal{A})_{\mathcal{M}}=\{0\}$), the mapping $G$ (respectively, the mapping $H$) is linear.
@@
$(i)$ We shall show that, for every $a\in\mathcal{A}$, the mapping $\Psi_{a}$ is linear. Given $a,b,c\in\mathcal{A}$ and $\lambda\in\mathbb{C}$, we have
@@
Hence, $\Psi_{a}$ is a linear mapping for any $a\in\mathcal{A}$. Similarly, one can show that the mapping $\Gamma_{a}:\mathcal{A}\rightarrow\mathcal{M},$ $\Gamma_{a}(t)=G(t)a$ is linear.
@@
$H(\lambda(b+c))=\lambda H(b)+\lambda H(c),$
for all $b,c\in\mathcal{A}$ and all $\lambda\in\mathbb{C}$, and thus $H(a+b)=H(a)+H(b),$ for all $a,b\in\mathcal{A}$, $H(0)=0$ and $H(\lambda a)=\lambda H(a)$ for all $\lambda\in\mathbb{C}$. The other statement can be similarly obtained. ∎
@@
Let $\mathbb{Z}$ be the set of all integers. Set $\mathcal{A}=\left\{\left[\begin{array}[]{cc}2n&0\\ 0&2n\end{array}\right]\ :\ n\in\mathbb{Z}\right\}.$ It is clear that $\mathcal{A}$ is a non-unital ring. Let =$\mathcal{M}=\left\{\left[\begin{array}[]{cc}i&j\\ 0&k\end{array}\right]\ :\ i,j,k\in\mathbb{Z}\right\}.$ A straightforward verification shows that $lann(\mathcal{A})_{\mathcal{M}}=\{0\}=rann(\mathcal{A})_{\mathcal{M}}$.
If $\mathcal{A}$ is unital, $lann(\mathcal{A})=\{0\}=rann(\mathcal{A})$. Also, if $\mathcal{A}$ is a semiprime algebra, then it is clear that $lann(\mathcal{A})=rann(\mathcal{A})=\{0\}$. If $\mathcal{A}$ satisfies Condition (P) in [12], then it is routine to see that $lann(\mathcal{A})=rann(\mathcal{A})=\{0\}$; where an algebra $\mathcal{A}$ satisfies Condition (P) if $aa_{0}a=\{0\}$ for any $a\in\mathcal{A}$ implies that $a_{0}=0$.
The dual space, $\mathcal{A}^{*}$, of a Banach algebra $\mathcal{A}$ is a Banach $\mathcal{A}$-bimodule with module operations given by $(\varphi a)(b)=\varphi(ab)$ and $(a\varphi)(b)=\varphi(ba)$, for all $a,b\in\mathcal{A}$, $\varphi\in\mathcal{A}^{*}.$ If $\mathcal{A}$ satisfies a Cohen factorization type property (i.e. for every $c\in\mathcal{A}$ there exists $a,b\in\mathcal{A}$ with $c=ab$), we have $lann_{\mathcal{A}^{*}}(\mathcal{A})=\{0\}=rann_{\mathcal{A}^{*}}(\mathcal{A})$. Recall that, by Cohen’s factorization theorem [20, Corollary 2.26], every Banach algebra with a bounded left approximate unit satisfies such a factorization property. To see the statement concerning modular annihilators, if $\varphi\in lann_{\mathcal{A}^{*}}(\mathcal{A}),$ we have $\varphi(ab)=0$ for all $a,b\in\mathcal{A}$, and thus, the factorization property implies that $\varphi=0$.
Our next proposition shows that every continuous generalized derivation of the third type from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule is a generalized derivation of the first and second type.
@@
Let $\mathcal{A}$ be a C-algebra, and let $F:\mathcal{A}\to\mathcal{M}$ be a continuous generalized derivation of the third type from $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule. Then $F$ is a generalized derivation of the first and second type.
@@
It is known that the product of $\mathcal{A}$ and the module products on $\mathcal{M}$ can be extended to a product on $\mathcal{A}^{**}$ and $\mathcal{A}^{**}$-bimodule operations on $\mathcal{M}^{**}$ via the first (or the second) Arens extensions [10, Theorem 2.6.15$(iii)$], respectively. Since $\mathcal{A}$ is a C-algebra, its product is Arens regular, that is, the first and second Arens products on $\mathcal{A}^{**}$ coincide, and the latter is a von Neumann algebra with respect to this product (cf. [10, Corollary 3.2.37]). It is also known that the following properties holds: for each $a\in\mathcal{A}$, $\tilde{a}\in\mathcal{A}^{**}$, $x\in\mathcal{M}$ and $z\in\mathcal{M}^{**}$, the mappings $y\mapsto ay$, $y\mapsto y\tilde{a}$ (respectively, $b\mapsto xb$, $b\mapsto bz$) are weak continuous maps on $\mathcal{M}^{**}$ (respectively, from $\mathcal{A}^{**}$ to $\mathcal{M}^{**}$) [10, Proposition A.3.52]; if $(a_{\lambda})$ and $(x_{\mu})$ are nets in $\mathcal{A}$ and $\mathcal{M}$, respectively, such that $a_{\lambda}\to a\in\mathcal{A}^{**}$ in the weak topology of $\mathcal{A}^{**},$ and $x_{\mu}\to x\in\mathcal{M}^{**}$ in the weak topology of $\mathcal{M}^{**}$, then
$ax=\lim_{\lambda}\lim_{\mu}a_{\lambda}x_{\mu}\hbox{ and }xa=\lim_{\mu}\lim_{\lambda}x_{\mu}a_{\lambda},$ (2.2)
in the weak topology of $\mathcal{M}^{**}$ (cf. [10, (2.6.26)]).
@@
and hence the uniform boundedness principle assures the existence of a positive $K_{1}$ satisfying $\|G(\cdot)b\|\leq K_{1}$ for all $b\in\mathcal{A}$ with $\|b\|\leq 1$. Similarly, there exists a positive $K_{2}$ satisfying $\|aH(\cdot)\|\leq K_{2}$ for all $a\in\mathcal{A}$ with $\|a\|\leq 1$.
Let us take an approximate unit $(u_{j})_{j}$ in $\mathcal{A}$. If we fix an element $a\in\mathcal{A},$ the net $(au_{j})_{j}$ converges in norm to $a$, and hence, by the continuity of $F$, $F(au_{j})$ tends to $F(a)$ in norm. It is also know that $(u_{j})_{j}\to\mathbf{1}$ in the weak-topology of $\mathcal{A}^{**}$, where $\mathbf{1}$ stands for the unit in $\mathcal{A}^{**}$. We therefore deduce from (2.2) that $(G(a)u_{j})_{j}\to G(a)\mathbf{1}$ in the weak-topology of $\mathcal{M}^{**}$. Thus, the identity $F(au_{j})=G(a)u_{j}+aH(u_{j})$ implies that
@@
It is easy to check that the mapping $R:\mathcal{A}\to\mathcal{M}^{**},$ $a\mapsto R(a)$ is linear. Moreover, since by the properties shown above $\|aH(u_{j})\|\leq K_{2}\|a\|$ for all $j$, we obtain $\|R(a)\|\leq K_{2}\|a\|$ for all $a\in\mathcal{A}$, and thus $R$ is continuous.
@@
$R(ab)=w^{*}\hbox{-}\lim_{j}abH(u_{j})=aw^{*}\hbox{-}\lim_{j}bH(u_{j})=aR(b)\ \ (\forall a,b\in\mathcal{A}),$
which guarantees that $R$ is a right multiplier from $\mathcal{A}$ to $\mathcal{M}^{**}$. Furthermore, by what we proved above, $\|R(u_{j})\|\leq K_{2}$ for all $j$, and hence, by the weak-compactness of the closed unit ball of $\mathcal{M}^{**}$, there exists a subnet, denoted again by $(u_{j})_{j}$, such that $\lim_{j}R(u_{j})=\xi\in\mathcal{M}^{**}$ in the weak-topology of $\mathcal{M}^{**}$. Since $R(au_{j})=aR(u_{j})$ for all $j$, and $(au_{j})_{j}\to a$ in norm, the continuity properties of the module operations on $\mathcal{M}^{**}$ and of the mapping $R$ give $R(a)=a\xi$ for all $a\in\mathcal{A}$.
@@
We shall next state two classical arguments in results on automatic continuity which are considered here from a more general point of view.
@@
Let $Z$ be a closed subspace of a Banach space $X$ such that $X/Z$ is finite-dimensional, and let $Y$ be a normed space. Suppose $F:X\to Y$ is a linear mapping whose restriction to $Z$ is continuous. Then $F$ is continuous.
@@
Since the quotient $X/Z$ is finite-dimensional, the subspace $Z$ is topologically complemented in $X$, that is, there exists a continuous linear projection $P:X\to X$ whose image is $Z$ and $Z^{\prime}=(Id-P)(X)$ is finite-dimensional (cf. [40, Lemma 4.21]). Clearly, by the finite-dimensionality of $Z^{\prime}$, $F|_{Z^{\prime}}:Z^{\prime}\to Y$ is continuous. Since $F|_{Z}$ is continuous by hypothesis, and $F(x)=F(P(x))+F((Id-P)(x))=F|_{Z}(P(x))+F|_{Z^{\prime}}((Id-P)(x)),$ for all $x\in X$, the conclusion is clear. ∎
@@
The desired conclusion is clear when $\mathcal{A}$ is unital, since in that case $F(a)=FL_{\mathbf{1}}(a)$. In the general case, we observe that the bilinear mapping $(a,b)\mapsto V(a,b):=F(ab)$ is separately continuous by hypothesis, so by the uniform boundedness principle, $V$ is jointly continuous. Thus, there exists a positive $K$ such that $\|F(ab)\|\leq K\ \|a\|\ \|b\|$ for all $a,b\in\mathcal{A}$. For each positive element $c\in\mathcal{A}$ with $\|c\|\leq 1$ there exists a positive $d$ with $d^{2}=c$, $\|d\|^{2}=\|d^{2}\|=\|c\|\leq 1$. Therefore $\|F(c)\|=\|F(d^{2})\|\leq K\|d\|^{2}=K\|c\|\leq K$. In particular, $F$ is bounded on the closed unit ball of $\mathcal{A}$. ∎
A cornerstone result in the theory of C-algebras, obtained by Cuntz in [9], asserts that a semi-norm $p$ on a C-algebra $\mathcal{A}$ which is bounded on each commutative self-adjoint subalgebra of $\mathcal{A}$, is bounded on the whole of $\mathcal{A}$. Cuntz’ theorem has been employed in results on automatic continuity of derivations, for example, in [38] Russo and the second author of this note apply it to prove that every Jordan derivation from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule is continuous and an associative derivation. More recently, An and He employ Cuntz’ theorem to prove that if $n\neq m$, then zero is the only $(m,n)$-Jordan derivation from a C-algebra into a Banach bimodule (cf. [3]). Next, we give another application by adapting the arguments in the proof of Theorem 2.8.
We are now ready to prove one of the main results of this section, which confirms that every generalized derivation of the third type from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule is continuous.
@@
Let $\mathcal{A}$ be a C-algebra, and let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule. Suppose that $F:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the third type. Then $F$ is continuous. Consequently, every generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the second type.